OpenStudy (anonymous):

Use elimination to solve. 2x+5y=9 2x+y=13

OpenStudy (jim766):

to eliminate a variable, look for ways to create opposites If you multiply -1 to the 1st equation that first 2x will be come -2x when added to 2x it will be eliminated -1(2x + 5y = 9) -2x -5y = -9 2x + y = 13 ------------ now add them both straight down

OpenStudy (anonymous):

The answer is (7,-1) ?